The relative permittivity of a dielectric is 2.25



A) what is its permittivity?



B) what will be the capacitance if this inserted into a 7.8 uF air capacitor




1
Expert's answer
2022-03-06T18:06:37-0500

"\varepsilon_m=\varepsilon_r\varepsilon_0=19.9\cdot 10^{-12}~\frac Fm,"

"C'=\frac{C}{\varepsilon_r}=3.5~\mu F."